OpenStudy (anonymous):

The sine of an angle whose terminal side falls in the first quadrant is .4067. Find the cosine of this angle. (Hint: Use the Pythagorean identity.) a. .165 b. .9136 c. .4067 d. not enough information is given

OpenStudy (mertsj):



OpenStudy (mertsj):

Use the Pythagorean Theorem. Or maybe just remember the Pythagorean Identity: \[\sin ^x+\cos ^2x = 1\]

OpenStudy (mertsj):

\[\sin ^2x+\cos ^2x=1\]

OpenStudy (mertsj):

Sorry for the previous typo

OpenStudy (anonymous):

using the Pythagorean identity: \[\sin ^{2}\theta + \cos ^{\theta} = 1\] \[\cos ^{2}\theta = 1-(0.4067)^{2} \] now just solve for cos

OpenStudy (anonymous):

I'm still not coming up with one of those answers.

OpenStudy (anonymous):

what do you get when you take 1-(0.4067)^2 ?

OpenStudy (mertsj):

Did you remember to hit the square root button after you did 1 - .4067 squared?

OpenStudy (anonymous):

that expression will \[=\cos ^{2}\theta \]

OpenStudy (anonymous):

@jonnymiller, I got .83459511

OpenStudy (anonymous):

then just take the square root of that

OpenStudy (anonymous):

I got θ=2 π ℕ_1 + arccos(((3 sqrt(9273279))/(10000)))

OpenStudy (anonymous):

i think you are making it more complicated....... set it up again, \[\cos ^{2}\theta + \sin ^{2}\theta = 1\] \[\cos ^{2}\theta + (0.4067)^{2} = 1 \] \[\cos ^{2}\theta = 1-0.16540489\] \[\cos ^{2}\theta =0.83459511\] \[\cos \theta = 0.9135617713\]

OpenStudy (anonymous):

Oh, okay. That makes sense. I was making it way too complicated, haha. Thank you for explaining!

OpenStudy (anonymous):

no problem, i hope you understand

OpenStudy (anonymous):

I do, thank you.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

:)